1. A shoes manufacturing company finds from the sales data of the other manufactures of similar product that the demand functions for the shoes can be expressed as
Q=10,000-20 P
Find:-
a. Number of shoes sold when P=$300
b. Price for selling 5400 pair of shoes
c. Price of zero sales
d. Point elasticity of demand at price $500
1
Expert's answer
2020-12-15T11:24:55-0500

Solution:

a.). Number of shoes sold when P = $300

Substitute for P in the demand function:

"Q =10,000 -(20\times 300)"

"Q=10,000-6,000"

"Q=4000"


Number of shoes sold = 4000 shoes


b.). Price for selling 5400 pair of shoes:

Substitute for Q in the demand function:

"5400 =10,000-20P"

"20P =10,000-5400"

"20P = 4,600"

"P= \frac{4,600}{20} = 230"


The price for selling 5400 pair of shoes is $230


c.). Price of zero sales:

Substitute for Q in the demand function:

"0=10,000-20P"

"20P=10,000"


"P=\frac{10,000}{20} = 500"


The price of zero sales is $500


d.). Point elasticity of demand at price $500:

From the demand function: Q=10,000 - 20P, the

"PED=\frac{∆Q}{∆P} \times \frac{P}{Q}"


"\frac{∆Q}{∆P}= -20"


"PED=(-20) \times (\frac{50}{0} ) = 0"


Point Elasticity of Demand = 0